OpenStudy (ms_keshaah):

With a wooden ruler you measure the length of a rectangular piece of sheet metal to be 12 \({\rm mm}\) . You use micrometer calipers to measure the width of the rectangle and obtain the value 5.98 \({\rm mm}\) . Give your answers to the following questions to the correct number of significant figures. a.What is the area of the rectangle? b.What is the perimeter of the rectangle? c.What is the difference between the length and width?

OpenStudy (sapphiremoon):

a) The area of the rectangle is length x width = 12 x 5.98 = 71.76 mm, which rounds to 72 mm because of significant figures (in multiplication, you round to the number of sig figs that the number with the least of them has). b)The perimeter of the rectangle is 2(length) + 2(width) = 2(12) + 2(5.78) = 24 + 11.56 = 35.56 which rounds to 36mm. (in addition, round to least number of decimal places) c)The difference between the length and width is 12 - 5.98 = 6.02 which rounds to 6mm. Hope this helps.